OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x - 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

\[(a + b)(a - b) = a^2 - b^2\] (\[(a + b)(a + b) = (a+b)^2 = a^2 + b^2 + 2ab\] Use these formulae for finding the values..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome dear..